OpenStudy (anonymous):

lim of [sin(pie+∆x)]/∆x as ∆x→0

OpenStudy (anonymous):

so basically you are trying to find the derivative of the function sin(x)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

that isn't even the definition of the derivative

OpenStudy (jamesj):

Actually it is. Recall first that the derivative of f at a is defined as the limit as h -> 0 of (f(x+h) - f(x))/h Here f = sin, and x = pi The ratio is (sin(pi+h) - sin(pi))/h = sin(pi+h)/h, as sin(pi) = 0. So the limit of this ratio as h -> 0 is d(sin x)/dx evaluated at x = 0. Now, the derivative of sin(x) is cos (x). Hence your limit is equal to ...

OpenStudy (anonymous):

he's missing -f(c)

OpenStudy (anonymous):

\[\lim _{\Delta x \rightarrow0} \sin(\pi+\Delta x) / \Delta x\]

OpenStudy (anonymous):

if he's doing the definition it should be\[\frac{f(c+\Delta x)-f(c)}{\Delta x}\]

OpenStudy (jamesj):

@willking: yes @outkast: no, because sin(pi) = 0.

OpenStudy (jamesj):

In your notation f(c) = sin(pi) = 0.

OpenStudy (anonymous):

ohh but pshh the question should actually state that... it shouldn't just totally detract the statement from the definition

OpenStudy (jamesj):

So ... @willking9 ... go back to my main answer. Do you follow and can you now see the answer?

OpenStudy (anonymous):

also you cannot assume that d/dx[sin]=cos

OpenStudy (anonymous):

If i were asked to use the definition and just wrote sin(x)=cos(x) .... i'd get it wrong

OpenStudy (anonymous):

You need a trig identity

OpenStudy (jamesj):

How do you know you can't assume the derivative of sin? But let's assume you're right. Show us how smart you are by showing us the trig identity in action: Simplify and evaluate the difference quotient (sin(x+h) - sin(x))/h as h --> 0

OpenStudy (anonymous):

Because the point of using the definition of the derivative is to find out the derivative Without an assumption

OpenStudy (jamesj):

Ok ... so, show us how it's done, enough postering. What have you got?

OpenStudy (jamesj):

I'll do it then. sin a - sin b = 2 sin( (a-b)/2 ) . cos( (a+b)/2 ) So ( sin(pi + h) - sin(pi) ) / h = 2 sin(h/2) . cos (pi + h) / h = cos(pi+h) . [ sin(h/2)/(h/2) ] Now, it is case that the limit as x -> 0 of sin(x)/x = 1, so this last term in square brackets, as h -> 0, is 1, while the limit of cos(pi+h) is cos(pi) Hence the limit as h -> 0 of ( sin(pi + h) - sin(pi) ) / h = cos(pi).1 = -1 **************

OpenStudy (anonymous):

or like i was saying do the proof of d/dx[sin(x)]=cos(x) and evaluating at c in which you would use the Trig Identity sin(c+h)=sin(c)cos(h)+cos(c)sin(h)

OpenStudy (zarkon):

Just use the fact that \[\sin(\pi+\Delta x)=-\sin(\Delta x)\] then \[\lim _{\Delta x \rightarrow0} \frac{\sin(\pi+\Delta x)}{\Delta x}=\lim _{\Delta x \rightarrow0} \frac{-\sin(\Delta x)}{\Delta x}=-1\]

OpenStudy (jamesj):

Ah, that's elegant.

OpenStudy (anonymous):

i was told to use the trig identity that states that sin(A+B)=sinAcosB+cosAsinB

OpenStudy (zarkon):

\[\sin(\pi+\Delta x)=\sin(\pi)\cos(\Delta x)+\cos(\pi)\sin(\Delta x)\] \[=0\cos(\Delta x)+(-1)\sin(\Delta x)=-\sin(\Delta x)\]